OpenStudy (anonymous):

What is the distance between point (2, -2) and (5,3) A. 3.2 units B. 4.3 units C. 5.8 units D. 7.1 units

OpenStudy (dtan5457):

Plug it into the distance formula.

jimthompson5910 (jim_thompson5910):

@vicyoria15, do you know how to use the distance formula dtan5457 posted?

OpenStudy (anonymous):

no not really @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what you do is basically subtract the corresponding coordinates and square that difference so for the x, we subtract: 2-5 = -3 then square -3 to get +9 for the y coordinates, what do you get when you subtract and square the result?

OpenStudy (anonymous):

-2-3 is -5 and the you square it and get...25?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

now you add those squares: 9+25 = 34 the last step is to take the square root of that to get \(\Large \sqrt{34}\)

OpenStudy (anonymous):

which is 5.8?

jimthompson5910 (jim_thompson5910):

that's the exact distance to find the approximate distance, you would use a calculator to evaluate \(\Large \sqrt{34}\)

jimthompson5910 (jim_thompson5910):

correct, roughly 5.8

OpenStudy (anonymous):

Thank you SOOO much

jimthompson5910 (jim_thompson5910):

so all that was using the distance formula